OpenStudy (anonymous):

what is 8x^2 = 72, how do you solve it

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

8x^2=72 x^2=72/8 x^2=9 x=3